OpenStudy (anonymous):

2(4x + 3) - 9 = 45 x = __________

OpenStudy (uri):

8x+6x -9 =45 14 x =45 +9 14 x=54 x =54/14 Divide and get the answer :D

OpenStudy (anonymous):

3.857142857

OpenStudy (uri):

Yes :D

OpenStudy (anonymous):

what would be the answer 3.8 or 3.9

OpenStudy (uri):

3.8?

OpenStudy (anonymous):

do i round the decimal

OpenStudy (uri):

No just write the answer which came :P